OpenStudy (barrelracing):

What are the discontinuities of the function f(x) = the quantity of x squared plus 5 x plus 6, all over 2 x plus 16.? a) x= -3 b) x = -2 c) x =-8 d) x = -5

OpenStudy (anonymous):

Do you have a calculator?

OpenStudy (barrelracing):

yes

OpenStudy (anonymous):

Take each of the choices and plug it in for x

OpenStudy (campbell_st):

so the problem is \[\frac{x^2 + 5x + 6}{2x + 16}\] so you just need to solve the denominator 2x + 16 = 0 you can't divide anything by zero so you need to find the value of x that makes the denominator zero that is the discontinuity

OpenStudy (barrelracing):

im confused why would i do that?

OpenStudy (campbell_st):

ok... you said you have a calculator what is the value of \[10 \div 0 = ?\]

OpenStudy (barrelracing):

there is no value

OpenStudy (campbell_st):

if you did the calculation you probably got a response like Math Error which means an answer doesn't exist... so a discontinuity is where part of the curve doesn't exist. so to find the discontinuity in this question you need to find the value of x that makes the denominator zero....

OpenStudy (campbell_st):

here is a graph of the cruve... the vertical dotted line is called an asymptote, it is a discontinuity... as you may see the curve doesn't cross the line. the value of the discontinuity is the the value of x that makes the denominator zero

OpenStudy (campbell_st):

so as I said earlier to find the discontinuity you need to solve 2x + 16 = 0 and that value of x is the discontinuity or you could plug all 4 choices into a calculator and then find the one that gives math error which is easier..?

OpenStudy (barrelracing):

ok